njjnjnjnjnj:

The temperature in Nico’s town was Negative 2.8 degrees Fahrenheit at noon. The temperature increased at a steady rate of mc011-2.jpg per hour until 6:00 p.m. Then, the temperature decreased by a total of 3.7 degrees Fahrenheit by midnight. What was the temperature at midnight?

leila123456:

k what do you think the answer is

justjm:

Work step-by-step. Translate the words into math. "The temperature in Nico’s town was Negative 2.8 degrees Fahrenheit at noon." \[-2.8°F\] ". The temperature increased at a steady rate of mc011-2.jpg per hour until 6:00 p.m." Can't see the rate because you forgot to attach the picture. If you come online then kindly do attach it with the "Attach File" or drag it from your desktop to the textbox. For now, I will denote the steady rate as \(r\) in units °F/hr. \[-2.8°F + (r \frac{°F}{hr})(6~hr) \] "Then, the temperature decreased by a total of 3.7 degrees Fahrenheit by midnight." \[-2.8°F + (r \frac{°F}{hr})(6~hr) - 3.7°F \] Therefore, the temperature at midnight is \(-2.8°F + (r \frac{°F}{hr})(6~hr) - 3.7°F \)

justjm:

Did a bit of research and the steady rate that is supposedly in the picture is 1.3 °F/hr. All you'll do is replace r with that rate.